Determine the empirical and molecular formula given the percent compositions or masses.


8.81 g Carbon, 91.2 g Chlorine, Molar mass: 1362.5 g/mol

1
Expert's answer
2021-11-19T00:00:01-0500

C"=12.011" Cl"=35.453"


"8.81g" Carbon "=\frac{8.81}{12.011}=0.7335moles"



"91.2g" Chlorine "=\frac{91.2}{35.453}=2.5724moles"



Mole ratio C: Cl "=0.7335:2.5724"

"=1:3.5"

"=2:7"


C2 Cl7 has a mass of "2(12.011)+7(35.453)"

"=272.193"



"\frac{1362.5}{272.193}=5.00"


Molecular Formulae "= _5("C2 Cl7 ")"

"=" C"_{10}" Cl"_{35}"